deduce ohm’s law by elementary ideas and hence write an expression for resistance and resistivity

Grade:12

1 Answers

Umakant biswal
7 years ago
dear malavika 
the elementary ideas of ohms law states that potential diffrence across a conducter is directly proportional to current passing through the conducter when physical condition will not change . 
so, v is directly prop to i 
and v/i = constant ( R ) 
where r is constant and known as resistance of the conducter , resistance is directly prop to the length and inversely prop , to the area of the cross section . 
therefore , 
R is directly prop to l/a 
on removing the sign of proportionality we get 
R = rho *l /a , where rho is the resistivity 
R will depend on nature of conducter , 
now , I = neAvd (vd is drift velocity ) 
and n is no of electrons present in unit volume of conducter . 
so, vd = eET/ m 
putting the value of vd in abvove equation we get 
I = neA * eET/m
further , we know that 
E = v/l so, I= neA * evt/ml 
again , V/I = R = ml/nesquare a *t 
and rho = R *a/l = m / n esquare t 
these are the required expression 
HOPE IT CLEARS 
REGARDS